# Prosoluble subgroups of the profinite completion of the fundamental group of compact 33 33-manifolds

### v2

@@ 4 Closed subgroups of fundamental groups of finite graphs of profinite groups with finite edge groups
In this section we will conclude an important step to prove Theorem E. The main theorem of this section give us necessary conditions to a prosoluble group be a subgroup of the fundamental group of an injective $k$-acylindrical finite graph of profinite groups with finite edge groups.
@@
Let $G$ be the fundamental group of an injective finite graph of pro-$\mathcal{C}$ groups. Suppose that there is an edge $e$ such that $G(e)=1$ and $G(d_{i}(e))\neq 1$ for $i=0,1$. Then $G$ splits as a free profinite product.
@@
[…]
be a slice of the graph $\Gamma$. Suppose that after remove the edge $e$ the graph $\Gamma$ is not connected:
[…]
Let $\Gamma_{1},\Gamma_{2}$ be the connected components and $G_{1},G_{2}$ its fundamental groups, respectively. They are embedded in $G$ by Lemma 2.11. It follows that $G=G_{1}\amalg G_{2}$. The other case is similar: if removing $e$ the graph remains connected
[…]
then we can choose a maximal subtree $T$ not containing $e$ and use it to calculate the fundamental group $G_{1}$ of the graph of groups restricted to $\Gamma-\{e\}$. So, $G$ will be a loop with edge correspondent to $G(e)$ and vertex correspondent to $G_{1}$ so that $G=\operatorname{HNN}(G_{1},G(e),t)=G_{1}\amalg\left\langle t\right\rangle$:
@@
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups with finite edge groups, $G$ its profinite fundamental group. Suppose $\mathcal{C}$ does not contain all finite groups. If $H$ is a pro-$\mathcal{C}$ subgroup of $G$, then there exists $e\in E(\Gamma)$ and a conjugate $H_{0}$ of $H$ such that $H_{0}\cap G(e)=1$ and for any such $e$
@@
where $U_{1}\simeq\pi_{1}(\mathcal{U},\Delta)$ is the profinite fundamental group of an injective $k$-acylindrical finite graph of profinite groups with finite edge groups, $\Delta$ a connected component of $\Gamma\setminus\{e\}$ and $U_{2}$ is either isomorphic to $\widehat{\mathbb{Z}}$ or similarly the profinite fundamental group of an injective $k$-acylindrical finite graph of profinite groups with finite edge groups.
@@
Consider the action of $H$ on the standard profinite tree $S=S(G)$. By Corollary 3.6 there is an edge $e$ of $S$ for which $H_{e}$ is trivial. Then there is an open subgroup $U_{H}$ of $G$ containing $H$ such that $(U_{H})_{e}=1$. By [ZM89, Corollary 4.5] $U_{H}$ is the profinite fundamental group of a finite graph of groups $(\mathcal{U},\Delta)$ whose vertex and edge groups are $U_{H}$-stabilizers of vertices and edges of $S$. Thus one of the edge groups $U(\bar{e})$ of $(\mathcal{U},\Delta)$ is trivial. Assuming, w.l.o.g., that $(\mathcal{U},\Delta)$ is reduced we can apply Proposition 4.1 to deduce that $U_{H}$ splits as a free profinite product. As $U(\bar{e})$ is conjugate by some element $G$ in $G$ into some edge group of $(\mathcal{G},\Gamma)$ we may conjugate $H$ and $U$ by this elelment to get the required $H_{0}$. This finishes the proof. ∎
@@
Let $G=\Pi_{1}(\mathcal{G},\Gamma)$ be the fundamental profinite group of a $k$-acylindrical finite graph of profinite groups. Let $v,w$ be vertices at distance $\geq 2k+1$. Then $\left\langle G(v),G(w)\right\rangle=G(v)\amalg G(w)$.
@@
Let $G=\Pi_{1}(\mathcal{G},\Gamma)$ be the profinite fundamental group of a finite $k$-acylindrical graph of profinite groups. Let $G(v)$, $G(w)$ be vertex groups of vertices at distance $\geq 2k+1$ having epimorphisms $f_{v}:G(v)\to C_{p}$, $f_{w}:G(w)\to C_{q}$ to cyclic subgroups of $S_{n}$ of order $p\neq q$ respectively. Then $f_{v},f_{w}$ extend to homomorphism $f_{vw}:\left\langle G(v),G(w)\right\rangle\to S_{n}$ that in turn extends to a homomorphism $f:U\to S_{n}$ from some open subgroup $U$ containing $\left\langle G(v),G(w)\right\rangle$.
Proof.
The existence of $f_{vw}$ follows from Proposition 4.5. The existence of $f$ then follows by [RZ10, Lemma 8.3.8]. ∎
Lemma 4.7.
Let $(\mathcal{G},\Gamma)$ be a $k$-acylindrical finite graph of profinite groups, $G$ its fundamental group and $S=S(G)$ its standard profinite tree. Suppose $\mathcal{C}$ does not contain all finite groups and $H$ is a pro-$\mathcal{C}$ subgroup of $G$. Then for $v,w\in V(S)$ such that $H_{v}\neq H_{w}$ are non-trivial and not both of order 22 22, there exists an open subgroup $U$ containing $H$ such that the natural images of $v$ and $w$ in $\Delta=U\backslash S$ are at distance $\geq 2k+1$.
Proof.
By Proposition 3.4, there is no edge $e$ such that
$|G_{d_{0}(e)}:G_{e}|=2=|G_{d_{1}(e)}:G_{e}|$
[…]
Theorem 4.8.
Let $(\mathcal{G},\Gamma)$ be a $k$-acylindrical finite graph of profinite groups and $G$ its fundamental group. Let $H$ be a subgroup of $G$ such that there exist some non-trivial non-conjugate subgroups $H\cap G(v)$, $H\cap G(w)$ which are not both pro-$p$ for any prime $p$. Then $H$ is not prosoluble.
Proof.
First observe that since $H\cap G(v)$, $H\cap G(w)$ are not both pro-$p$ we can replace $H$ by its open subgroup such that $H\cap G(v)$, $H\cap G(w)$ have finite cyclic quotients $C_{p}$ and $C_{q}$ for $p\neq q$, respectively.
By Lemma 4.7 there exists an open subgroup $U$ containing $H$ such that the natural images $\bar{v}$, $\bar{w}$ of $v$ and $w$ in $\Delta=U\backslash S$ are at distance $\geq 2k+1$. Moreover, for some such $U$ the vertex groups $U(v)$ and $U(w)$ have epimorphisms to finite cyclic quotients $C_{p}$ and $C_{q}$ for $p\neq q$, respectively which are non-trivial restricting to $H\cap G(v)$ and $H\cap G(w)$ (cf. [RZ10, Lemma 8.3.8]). Choose $C_{p},C_{q}\in S_{n}$, $n>4$, such that any conjugate of them generate non-soluble subgroup (see [Jar94, Lemmas 1.1 and 1.2]) and let $f_{v}:U(v)\to C_{p}<S_{n}$, $f_{w}:U(w)\to C_{q}<S_{n}$ corresponding epimorphisms. Then by Corollary 4.6 we can choose our $U$ such that $f_{v},f_{w}$ extend to an epimorphism $f:U\to S_{n}$. Then $f$ restricted on $H$ surjects on $S_{n}$ and so $H$ is not prosoluble. ∎
Corollary 4.9.
Let $(\mathcal{G},\Gamma)$ be a $k$-acylindrical finite graph of profinite groups and $G$ its fundamental group. Let $H$ be a prosoluble subgroup of $G$. Then one of the following holds:
@@
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups, $G$ its fundamental group, $S$ its standard graph and $H$ is a prosoluble subgroup of $G$. If $H$ is finitely generated, then it has only finitely many vertex stabilizers up to conjugation. In particular, if $D$ is the maximal abstract subgraph of $\Gamma$ such that each $e\in D$ satisfies $H_{e}\neq 1$, then $D$ has finitely many connected components up to translation.
@@
Consider the decomposition
@@
where $G_{U}=G/\tilde{U}$ with $\tilde{U}$ is the subgroup generated by the vertex stabilizers of $G$ and $G_{U}$ acts $k$-acylindrically on $S_{U}=\tilde{U}\backslash S$ by Theorem 3.8. Note that since $\tilde{U}$ is generated by all vertex stabilizers, then $G_{U}$ has finite stabilizers. Setting
$H_{U}=H\tilde{U}/\tilde{U}\simeq H/(\tilde{U}\cap H)$
we have the decomposition
[…]
It follows that $G_{U}$ and $H_{U}$ satisfies the hypothesis of Theorem 4.2. Besides, conjugating $H_{U}$ if necessary we may assume that $H_{U}=(H_{U})_{0}$ from Theorem 4.2 and therefore satisfies one of the itens (i), (ii) or (iii) of Corollary 4.3.
Suppose that $H_{U}$ satisfies Corollary 4.3 (i). Following the notation of Corollary 4.3, all intersections $H_{U}\cap uL_{i}u^{-1}$ are pro-$p$ for every $u$ and $L_{i}$ (we use the notation $L_{i}$ in the place of $U_{i}$ of Corollary 4.3 to avoid any confusion) and satisfies
$\sum_{i,u}d(H_{U}\cap uL_{i}u^{-1})\leq d(H_{U})\leq d(H).$
The groups $H_{U}\cap uL_{i}u^{-1}$ are finitely generated pro-$p$ groups whose induced action on $S_{U}$ is $k$-acylindrical. By [WZ18, Theorem 3.8] they decompose as a free pro-$p$ product of vertex stabilizers and a free pro-$p$ group where the number of free factors is bounded by $d(H_{U}\cap uL_{i}u^{-1})\leq d(H)$. Since the vertex stabilizers are finite, they are contained in some conjugate of $L_{i}$, that is, $(H_{U})_{v}\subset aL_{i}a^{-1}$ and the stabilizer of $v$ in $H_{U}\cap uL_{i}u^{-1}$ is
$(H_{U})_{v}\cap H_{U}\cap uL_{i}u^{-1}=(H_{U})_{v}.$
It follows that $H_{U}$ has finitely many vertex stabilizers up to conjugation for each $U$ and the number of this vertex stabilizers is bounded by the number of generators of $H$.
Now if $H_{U}$ satisfies (ii) then all vertex stabilizers are finite and conjugate (as they are Hall subgroups, cf. [RZ10, Theorem 2.3.5]).
Suppose $H_{U}$ satisfies Corollary 4.3 (iii) and $D_{U}=\{e\in S_{U}:(H_{U})_{e}\neq 1\}$. We show that $H\backslash D_{U}$ is connected. Suppose not. For any open $L$ as in Corollary 4.3 and let $D_{L}=\{e\in S_{U}\mid L_{e}\neq 1\}$. Then every connected component of $D_{U}$ is in some connected component of $D_{L}$ and $D_{U}=\bigcap_{L}D_{L}$. Hence there exists $L$ such that the number of connected components of $D_{L}$ is greater than 1 up to translation. This means the image of $D_{L}$ in $L\backslash S_{U}$ is disconnected and so $L=\coprod_{i}L_{i}\amalg F$ is a free profinite product of the fundamental groups $L_{i}=\Pi_{1}(\mathcal{L},\Delta_{i})$, where $\Delta_{i}$ are connected components of $L\backslash D_{L}$. Then one can write $L=L_{1}\amalg L_{2}$ with $L_{i}\leq L_{1}$, $L_{j}\leq L_{2}$ for some $i\neq j$. This means that there are $v\neq w\in S_{U}$ such that $H_{v}\leq U_{1}$ and $H_{w}\leq U_{2}$, contradicting the assumption Corollary 4.3 (iii). Thus $D_{U}$ has just one connected component up to translation in this case.
By Corollary 3.7 $H_{U}$ has a finite number of maximal vertex stabilizers up to conjugation not exceeding $d(H)$ and therefore the number of maximal vertex stabilizers of $H$ acting on $S$ does no exceed $d(H)$. Then the number of the connected components of $D$ up to translation does not exceed $d(H)$ as well. ∎

### v3

@@ 3 Acylindrical actions and prosoluble groups
Recall that for any open normal subgroup $U$ of a pro-$\mathcal{C}$ group $G$ acting on a pro-$\mathcal{C}$ tree $T$ we set
$\widetilde{U}=\left\langle U\cap G_{v}:v\in V(T)\right\rangle,$
and $G_{U}=G/\widetilde{U}$, $\T\displaystyle T_{U}=\widetilde{U}\backslash$ T.
Proposition 3.9.
Suppose $\mathcal{C}$ does not contain all finite groups and let $G$ be a pro-$\mathcal{C}$ group acting $k$-acylindrically on a simply connected profinite tree $T$. The group $G_{U}$ acts on a profinite simply connected $G_{U}$-quotient tree $T^{\ast}_{U}$ of $T_{U}$ with trivial edge stabilizers and vertex stabilizers equal to the maximal (by inclusion) stabilizers of vertices of $T_{U}$. Moreover, if $G$ is prosoluble then $(G_{U},V(T^{\ast}_{U}))$ is $\mathcal{S}$-projective, where $\mathcal{S}$ is the class of all finite soluble groups.
Proof.
By Proposition 2.19 $D_{U}=\{m\in\widetilde{U}\backslash T\mid(G/\widetilde{U})_{m}\neq 1\}$ is a profinite subgraph of $T_{U}$. We can collapse all connected components of the profinite subgraph $D_{U}$ of $T_{U}$ (see [Rib17, Exercise 2.1.11]); the resulting profinite quotient graph $T^{\ast}_{U}$ of $T_{U}$ is $G$-invariant and by [Rib17, Proposition 3.9.1] is simply connected. The edge stabilizers of $T^{\ast}_{U}$ are trivial; the vertices of $T^{\ast}_{U}$ whose stabilizers are non-trivial are exactly the vertices to which the connected components of $D_{U}$ were collapsed; thus the only non-trivial vertex stabilizers of $T^{\ast}_{U}$ are stabilizers of the connected component $\operatorname{Stab}(D_{U})_{\ast}$ of $D_{U}$. But as was explained in the proof of Theorem 3.8 any connected component of $D_{U}$ has diameter $\leq 2k$. Hence $\operatorname{Stab}(D_{U})_{\ast}$ stabilizers a vertex in $(D_{U})_{\ast}$ (cf. [Ser02, Corollary, Page 20]) and clearly the stabilizer of this vertex is maximal by inclusion. This concludes the first part of the proof.
Suppose now $G$ is prosoluble. By item (i) of Theorem 2.24 $(H_{U},V(T^{*}_{U}))$ is $\mathcal{S}$-projective that finishes the proof of second statement. ∎
Lemma 3.10.
Let $G$ be a prosoluble group acting $k$-acylindrically on a simply connected profinite tree $T$. Suppose for some primes $p\neq q$ and some $v,w\in V(T)$ from different $G$-orbits, the stabilizer $G_{v}$ has a subgroup of order $p$ and the stabilizer $G_{w}$ has a subgroup of order $q$ as subquotients. Then there exists a subgroup $H$ of $G$ that admits an epimorphism $f:H\to C_{pq}$ to a cyclic group of order $pq$ such that $|f(H_{v})|=p,|f(H_{w})|=q$ and $K=\ker(f)$ is not generated by its vertex stabilizers. Moreover, non-trivial subgroups of the images of $H_{v}$ and $H_{w}$ in $H/\widetilde{K}$ do not commute up to conjugation.
Proof.
W.l.o.g we may assume that $G_{v}$ and $G_{w}$ are maximal (by inclusion) vertex stabilizers. Choose an open normal subgroup $U$ of $G$ such that $G_{v}U/U$ and $G_{w}U/U$ have subgroups $G_{p}$ of order $p$ and $G_{q}$ of order $q$, respectively. Consider the action of $G_{U}=G/\widetilde{U}$ on $\S\displaystyle T_{U}=\widetilde{U}\backslash$ S and note that in $G_{U}$ the vertex stabilizers of the images $v_{u},w_{u}$ of $v$ and $w$ in $T_{U}$ are $G_{v}U/U$ and $G_{w}U/U$ (still in different orbits) and refining $U$ we may assume that $G_{v}U/U$ and $G_{w}U/U$ are maximal vertex stabilizers. By Proposition 3.9 $G_{U}$ acts on a simply connected (infinite) profinite tree $T^{*}_{U}$ with trivial edge stabilizers, the vertex stabilizers of $T^{*}_{U}$ are maximal vertex stabilizers of $T_{U}$ and $(G_{U},V(T^{*}_{U}))$ is $\mathcal{S}$-projective, where $\mathcal{S}$ is the class of all finite soluble groups. Hence $G_{p}$ and $G_{q}$ are subgroups of some stabilizers of vertices of $T^{*}_{U}$ from distinct $G_{U}$-orbits. Note that if a subgroup of $G$ generated by its vertex stabilizers then its image in $G_{U}$ generated by its vertex stabilizers as well, and vertex stabilizers in $G_{U}$ are finite. Therefore to prove the lemma it suffices to construct a subgroup $H$ in $G_{U}$ containing $G_{p}$ and $G_{q}$ and an epimorphism $f:H\to C_{pq}$ to a cyclic group of order $pq$ with $|f(G_{p})|=p,|f(G_{q})|=q$ such that $\ker(f)$ is not generated by torsion.
Put $H=\langle G_{p},G_{q}\rangle$. Then by [Zal24, Lemma 4.8] setting
$\mathcal{H}=\{H\cap(G_{U})_{v}\mid v\in V(T_{U})\}$
there is a profinite $H$-space $X$ such that
  • (a)
    $\mathcal{H}=\{(H)_{x}\mid x\in X\}$;
  • (b)
    $(H)_{x}\cap(H)_{x^{\prime}}=1$ for $x\neq x^{\prime}$;
  • (c)
    $\mathbf{H}=(H,X)$ is a $\mathcal{C}$-projective pile.
Moreover, since $H$ is finitely generated, it is second countable and so by [Zal24, Remark 4.9] there exists a second countable such $X$. Therefore there exists a continuous section $\XX\displaystyle\sigma:H\backslash$ $\to X$ (see [RZ10, Lemma 5.6.7]).
Hence by Proposition 2.25 $H$ embeds into a free prosoluble product $H_{0}=\coprod_{s\in Im(\sigma)}H_{s}\amalg F$, where $F$ is free prosoluble, such that each $H_{x}=H_{s}^{h}$ for some $h\in H_{0}$. Choose $x_{1},x_{2}\in X$ such that $H_{x_{1}}$ contains $G_{p}$ and $H_{x_{2}}$ contains $G_{q}$, observe that $x_{1},x_{2}$ are from different $H$-orbits and let $s_{1}\neq s_{2}\in Im(\sigma)$ such that $H_{s_{1}}$ and $H_{s_{2}}$ conjugate to $H_{x_{1}}$ and $H_{x_{2}}$ respectively.
By [Rib17, Theorem 5.3.4] $H_{0}$ is an inverse limit of free prosoluble products with finitely many factors which are quotients free factors of $H_{0}$ and so there exists such a free prosoluble product $\coprod_{i=1}^{n}H_{s_{i}}\amalg F$ with finitely many factors, where $H_{s_{1}}$ and $H_{s_{2}}$ are still the free factors. Then factoring out all factors but $H_{s_{1}}$ and $H_{s_{2}}$ we arrive at a free prosoluble product $H_{s_{1}}\amalg H_{s_{2}}$, where $H_{s_{1}}$ and $H_{s_{2}}$ are still conjugate to the images of $H_{x_{1}}$ and $H_{x_{2}}$ respectively.
Now define a natural epimorphism $H_{s_{1}}\amalg H_{s_{2}}\to H_{s_{1}}\times H_{s_{2}}$ that sends $H_{s_{1}}$ onto $H_{s_{1}}$ and $H_{s_{2}}$ onto $H_{s_{2}}$ identically. The kernel of this natural epimorphism is free prosoluble (by [RZ10, Theorem 9.1.6]). Hence the kernel of $f$ factors through the kernel of the restriction of this epimorphism to the image $M_{U}$ of $H_{U}$ in $H_{s_{1}}\amalg H_{s_{2}}$ which is torsion-free. Therefore the kernel of $f$ is not generated by torsion as required. Moreover, as non-trivial subgroups of $H_{s_{1}},H_{s_{2}}$ in $H_{s_{1}}\amalg H_{s_{2}}$ do not commute up to conjugation (cf. [Rib17, Theorem 9.1.2]), hence non-trivial subgroups of $H_{x_{1}}$ and $H_{x_{2}}$ of $M_{U}$ do not commute up to conjugation as well. ∎
@@ 4 Closed subgroups of fundamental groups of finite graphs of profinite groups with finite edge groups
In this section we will conclude an important step to prove Theorem E. The main theorem of this section gives us necessary conditions for a prosoluble group to be a subgroup of the fundamental group of an injective $k$-acylindrical finite graph of profinite groups with finite edge groups. The fundamental groups of a finite graph of profinite groups are always profinite (or pro-$\mathcal{C}$) in this section unless otherwise said; if vertex groups are finitely generated they are just profinite completion of the classical Bass-Serre fundamental groups.
@@
Let $G$ be the fundamental pro-$\mathcal{C}$ group of an injective reduced finite graph of pro-$\mathcal{C}$ groups $(\mathcal{G},\Gamma)$. Suppose that there is an edge $e$ of $\Gamma$ such that $G(e)=1$ and $G(d_{i}(e))\neq 1$ for $i=0,1$. Then $G$ splits as a free $\mathcal{C} pro$duct of either fundamental groups of two subgraphs of groups of $(\mathcal{G},\Gamma)$ or as the fundamental group of the subgraph of groups of $(\mathcal{G},\Gamma)$ and $\widehat{\mathbb{Z}}_{\mathcal{C}}$.
@@
$G(e)\displaystyle\vdots\displaystyle\vdots$
be a slice of the graph $\Gamma$.
After removing the edge $e$ we have two possibilities: either the resulting graph is connected or not connected. We shall prove that $G$ splits as a free profinite product in each of the possibilities.
Case 1. After removing the edge $e$ the graph $\Gamma$ become not connected:
$\vdots\displaystyle\vdots$
Let $\Gamma_{1},\Gamma_{2}$ be the connected components and $G_{1},G_{2}$ its fundamental groups, respectively. They are embedded in $G$ by Lemma 2.11. It follows that $G=G_{1}\amalg G_{2}$.
Case 2. After removing $e$ the graph remains connected
$\vdots\displaystyle\vdots$
then we can choose a maximal subtree $T$ not containing $e$ and use it to calculate the fundamental group $G_{1}$ of the graph of groups restricted to $\Gamma-\{e\}$. So, $G$ is the fundamental group of the loop with edge correspondent to $G(e)$ and vertex correspondent to $G_{1}$ so that $G=\operatorname{HNN}(G_{1},G(e),t)=G_{1}\amalg\left\langle t\right\rangle$, since $G(e)=1$.
@@
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups with finite edge groups and $G$ its profinite fundamental group. Suppose $\mathcal{C}$ does not contain all finite groups. If $H$ is a pro-$\mathcal{C}$ subgroup of $G$, then there exists $e\in E(\Gamma)$ and a conjugate $H_{0}$ of $H$ such that $H_{0}\cap G(e)=1$ and for any such $e$
@@
where $U_{1}\simeq\Pi_{1}(\mathcal{U},\Delta)$ is the profinite fundamental group of an injective $k$-acylindrical finite graph of profinite groups with finite edge groups, $\Delta$ a connected component of $\Gamma\setminus\{e\}$ and $U_{2}$ is either isomorphic to $\widehat{\mathbb{Z}}$ or similarly is the profinite fundamental group of an injective $k$-acylindrical finite graph of profinite groups with finite edge groups.
@@
Consider the action of $H$ on the standard profinite tree $S=S(G)$. By Corollary 3.6 there is an edge $e$ of $S$ for which $H_{e}$ is trivial. Then there is an open subgroup $U_{H}$ of $G$ containing $H$ such that $(U_{H})_{e}=1$. By [ZM89a, Corollary 4.5] $U_{H}$ is the profinite fundamental group of a finite graph of groups $(\mathcal{U},\Delta)$ whose vertex and edge groups are $U_{H}$-stabilizers of vertices and edges of $S$. Thus one of the edge groups $U(\bar{e})$ of $(\mathcal{U},\Delta)$ is trivial. Assuming, w.l.o.g., that $(\mathcal{U},\Delta)$ is reduced we can apply Proposition 4.1 to deduce that $U_{H}$ splits as a free profinite product. As $U(\bar{e})$ is conjugate by some element $g\in G$ into some edge group of $(\mathcal{G},\Gamma)$ we may conjugate $H$ and $U$ by this element to get the required $H_{0}$. This finishes the proof. ∎
@@
Let $G=\Pi_{1}(\mathcal{G},\Gamma)$ be the profinite fundamental group of a $k$-acylindrical finite graph of profinite groups. Let $v,w\in V(\Gamma)$ be vertices at distance at least $2k+1$. Then $\left\langle G(v),G(w)\right\rangle=G(v)\amalg G(w)$.
@@
For the proof of the next theorem we shall need two simple lemmas.
Lemma 4.6.
Let $G$ be a pro-$\mathcal{C}$ group acting on a pro-$\mathcal{C}$ tree $T$. Let $\mathcal{U}=\{U_{i}\}$ be a filtered from below family of subgroups of $G$ with $H=\bigcap_{U\in\mathcal{U}}U$. Then $\widetilde{H}=\bigcap_{U\in\mathcal{U}}\widetilde{U}$.
Proof.
By [Rib17, Proposition 4.1.1] $\T\displaystyle\widetilde{U}\backslash$ T is a pro-$\mathcal{C}$ tree and so
[…]
is a pro-$\mathcal{C}$ tree (cf. [Rib17, Proof of Lemma 4.2.7]). Hence by [Rib17, Proof of Lemma 4.2.7] $\bigcap_{U\in\mathcal{U}}\widetilde{U}$ is generated by the vertex stabilizers and so equal to $\widetilde{H}$. ∎
Lemma 4.7.
Let $G$ be an extension of an abstract free group $F$ by a group $C_{pq}$ of order $pq$ for primes $p\neq q$. Suppose that $G$ contains finite subgroups $A$ and $B$ of order $p$ and $q$ that do not commute up to conjugation. Then the natural epimorphism $f:G\to C_{pq}$ with kernel $F$ factors through an abstract free product $C_{p}\ast C_{q}$.
Proof.
By the results of Karrass, Pietrowski, Solitar, Cohen and Scott [KPS73, Coh06, Sco74]) $G$ splits as the fundamental group $\Pi_{1}(\mathcal{G},\Gamma)$ of a finite graph of finite groups and assuming w.l.o.g. that $(\mathcal{G},\Gamma)$ is reduced, one sees that the vertex groups of $(\mathcal{G},\Gamma)$ are of order at most $pq$ and edge groups are of order at most $p$ or $q$.
As any finite group is conjugate into a vertex group, we shall assume w.l.o.g. that $A$ and $B$ are in the vertex groups of $(\mathcal{G},\Gamma)$. Let $(\mathcal{G},\Gamma_{A})$ be the maximal subgraph of groups of $(\mathcal{G},\Gamma)$ containing $A$ in its every vertex group and $(\mathcal{G},\Gamma_{B})$ is the maximal subgraph of groups of $(\mathcal{G},\Gamma)$ containing $B$ in its every vertex group. Since $A$ and $B$ do not commute up to conjugation, $\Gamma_{A}$ and $\Gamma_{B}$ do not intersect. Then factoring out the normal closure of all subgroups of order $q$ contained in the vertex groups of $(\mathcal{G},\Gamma_{A})$ and all subgroups of order $p$ contained in the vertex groups of $(\mathcal{G},\Gamma_{B})$ we obtain the fundamental group $\Pi_{1}(\mathcal{H},\Gamma)$, where $A$ is still contained in a vertex group of $(\mathcal{H},\Gamma_{A})$, $B$ is still contained in the vertex groups of $(\mathcal{H},\Gamma_{B})$ but neighboring edge groups with trivial edge groups of $\Gamma_{A}$ and $\Gamma_{B}$ are trivial. This means that $A$ and $B$ are free factors of $\Pi_{1}(\mathcal{H},\Gamma)$ and so so the group $\Pi_{1}(\mathcal{H},\Gamma)$ can be mapped to $A\ast B$. ∎
Theorem 4.8.
Let $(\mathcal{G},\Gamma)$ be a $k$-acylindrical finite graph of profinite groups, $G$ its profinite fundamental group and $S=S(G)$ its standard profinite tree. Let $H$ be a subgroup of $G$ such that there exist some non-trivial non-conjugate vertex stabilizers $H_{v}$, $H_{w}$ for $v,w\in S$ which are not both pro-$p$ for any prime $p$. Then $H$ is not prosoluble.
Proof.
[…]
Since $f_{H}(H_{v})$ has order $p$ and $f_{H}(H_{w})$ has order $q$, it follows from this commutative diagram that the image of $H$ in $C_{p}\amalg C_{q}$ contains a group of order $p$ and a group of order $q$.
Choose now elements $c_{p},c_{q}\in S_{n}$, $n>4$ of order $p$ and $q$, such that any conjugate of them generate non-soluble subgroups (see [Jar94, Lemmas 1.1 and 1.2]) and let $f_{p}:C_{p}\amalg C_{q}\to C_{p}<S_{n}$, $f_{q}:C_{p}\amalg C_{q}\to C_{q}<S_{n}$ be the corresponding epimorphisms. Then we have an epimorphism $\varphi:C_{p}\amalg C_{q}\to S_{n}$. Since the image of $H$ in $C_{p}\amalg C_{q}$ contains the group of order $p$ and the group of order $q$, it maps epimorphically to $S_{n}$ which is not soluble, a contradiction. This finishes the proof. ∎
@@
Let $(\mathcal{G},\Gamma)$ be a $k$-acylindrical finite graph of profinite groups and $G$ its profinite fundamental group. Let $H$ be a prosoluble subgroup of $G$. Then one of the following holds:
@@
Let $(\mathcal{G},\Gamma)$ be an injective $k$-acylindrical finite graph of profinite groups, $G$ its profinite fundamental group, $S$ its standard graph and $H$ is a prosoluble subgroup of $G$. If $H$ is finitely generated, then it has only finitely many maximal vertex stabilizers up to conjugation. In particular, if $D$ is the maximal abstract subgraph of $S$ such that each $e\in D$ satisfies $H_{e}\neq 1$, then $D$ has finitely many connected components up to translation.
@@
By Corollary 4.9 we have that $H$ satisfies (i), (ii) or (iii).
If $H$ satisfies (ii) and not (i) and (iii), then all vertex stabilizers are finite and conjugate (as they are Hall subgroups, cf. [RZ10, Theorem 2.3.5]).
@@
Suppose first that $H$ satisfies (i). Then by Proposition 3.9 and Theorem 3.8 $H_{U}$ acts acylindrically on a simply connected profinite tree $S^{\ast}_{U}$ with trivial edge stabilizers, vertex stabilizers equal to the maximal stabilizers of $S_{U}$ and $(H_{U},V(\mathcal{S}^{\ast}_{U}))$ is $\cal S$-projective. Moreover, by Theorem 3.5 the connected components of $D_{U}=\{e\cup d_{0}(e)\cup d_{1}(e)\mid(H_{U})_{e}\neq 1\}$ have finite diameter and so $S^{\ast}_{U}$ is non-trivial by Corollary 3.6. By [Zal24, Proposition 4.9], the number of non-conjugated $(H_{U})_{v_{\ast}}$, $v_{\ast}\in S^{\ast}_{U}$, is bounded by $d(H_{U})\leq d(H)$. Since the non-trivial vertex stabilizers of $S^{\ast}_{U}$ are stabilizers of connected components of $D_{U}$, it follows that the number of connected components of $D_{U}$ and hence the number of the maximal vertex stabilizers of $S_{U}$ is bounded by $d(H)$ up to conjugation, independently on $U$.
Suppose that $H$ satisfies (iii). If $H\leq G(v)^{g}$ then $H$ fixes a vertex and the same is true for each $H_{U}$. It follows that the number of maximal vertex stabilizers, up to conjugation, is trivially bounded by $d(H)$.
Thus, in both cases (i) and (iii) each $H_{U}$ has finitely many maximal vertex stabilizers, up to conjugation. This finishes the proof. ∎
Proof of Theorem G follows from Proposition 4.10.